Please show work and use a product of sines and cosines formula to rewrite cos(5t) sin(8t)

Answer by lwsshak3(11628)

Please show work and use a product of sines and cosines formula to rewrite cos(5t) sin(8t)
product-to-sum formula:
sin x cos y=1/2[sin(x+y)-sin(x-y)]
cos(5t) sin(8t)= sin(8t) cos(5t)
=1/2[sin(8t+5t)-sin(8t-5t)]
=1/2[sin(13t)-sin(3t)]